# Probing dark matter interactions with a RES-NOVA prototype cryogenic detector

### v1

@@ 4 Data Processing and detector performance
A simple event-selection criterion was adopted: the OF amplitude and the fit-derived amplitude were required to agree within 10%. This requirement ensures that only well-reconstructed pulses, whose shape is consistent with the expected detector response, are retained for the subsequent analysis.
@@
In Fig. 5, the background energy spectrum of all accepted events, including interactions and nuclear recoils, is shown. The spectrum exhibits a steep rise at low energies, around , which originates from the acceptance of events with a signal–to–noise ratio close to unity. The inclusion of these noise–dominated events is primarily driven by the limited signal sensitivity of the thermal sensors employed in this study, namely Ge–NTD thermistors, when compared to other sensor technologies better suited for low–energy dark matter searches, such as transition–edge sensors (TES) Pirro and Mauskopf (2017).
The dominant contribution to the total counting rate in the DM RoI (2.5–10 keV) originates from background sources external to the crystal. These are primarily associated with the cryogenic infrastructure, such as Bremsstrahlung radiation from 210Bi present in the internal Pb shielding, and with environmental backgrounds, mainly due to ambient -ray radiation, as previously discussed in Ref. Alloni and others (2025). To further illustrate this point, Fig. 5 also shows the expected contribution from radioactive bulk contaminations within the operated PbWO4 crystal. This component is evaluated through Monte Carlo simulations of the experimental setup, using as input the results of the high-statistics radioactive screening measurements performed on the same crystal and summarized in Tab. 1.
To evaluate the detection and reconstruction efficiency in the RoI for DM investigations, we injected a comb of simulated mono-energetic pulses directly into the raw data stream and processed them through the full analysis pipeline. For each simulated energy, events that passed the quality selection and were correctly reconstructed were counted as successful detections. The efficiency was then obtained as the fraction of successfully reconstructed events over the total number of injected pulses at each energy value. In Fig. 6, the results of this analysis step are shown. In addition, for the evaluation of the detector baseline resolution, we fit a Gaussian distribution to the simulated lines obtaining: 234 eV, see Fig. 7.
@@ Acknowledgements.
The RES–NOVA collaboration thanks the directors and staff of the Laboratori Nazionali del Gran Sasso for their support, as well as the technical staff involved in the experiment. This work received funding from the European Union’s Horizon Europe programme through the ERC grant ERC–101087295 RES–NOVA. Additional financial support was provided under the National Recovery and Resilience Plan (NRRP) by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU, within the project “Advanced techniques for a next–generation isotopically enriched cryogenic Dark Matter experiment” (2022L2AXP2). This work made use of the Arby software for Geant4–based Monte Carlo simulations, developed within the framework of the Milano–Bicocca activities and currently maintained by O. Cremonesi and S. Pozzi. We gratefully acknowledge the University of Milano–Bicocca (UNIMIB) and the Istituto Nazionale di Fisica Nucleare (INFN) for their continuous support of the collaboration. F.A. Danevich and V.I. Tretyak thank the Laboratori Nazionali del Gran Sasso, the Gran Sasso Science Institute, and the RES–NOVA Collaboration for their outstanding support and hospitality during the difficult period following the Russian invasion of Ukraine.References

### v2

@@ 4 Data Processing and detector performance
A simple event–selection criterion was adopted: the amplitude obtained from the OF was required to be consistent, within a given tolerance, with the amplitude extracted from the pulse fit. This condition ensures that only well–reconstructed pulses, whose shapes are compatible with the expected detector response, are retained for the subsequent analysis. Several tolerance values were tested independently, namely 5%, 10%, 20%, and 30%. Since this parameter strongly influences the final sample of accepted events, its variation provides an estimate of the systematic uncertainty associated with the selection procedure. For the analysis presented here we adopted a tolerance of 10%.
@@
In Fig. 5, the background energy spectrum of all accepted events, including interactions and nuclear recoils, is shown. The spectrum exhibits a steep rise at low energies, around , which originates from the acceptance of events with a signal–to–noise ratio close to unity. These events may include contributions of non-physical origin or signals whose pulse shape is inconsistent with the detector response model. Stricter selection criteria suppress the fraction of such noise–dominated events in the accepted sample, while looser criteria, conversely, lead to an increased acceptance of these events. At the same time, stricter criteria reduce the overall detection efficiency, whereas looser selections enhance it. This suppressing or enhancing effect is observable down to energies of about ; below this threshold, the impact of the selection criteria on the spectrum is no longer appreciable, indicating a signal–to–noise ratio below unity. The threshold for event identification is primarily determined by the signal sensitivity of the thermal sensors employed in this study, namely Ge–NTD thermistors. In comparison with other sensor technologies better suited for low–energy DM searches, such as transition edge sensors (TES) Pirro and Mauskopf (2017), significantly lower thresholds are expected.
The dominant contribution to the total counting rate in the DM RoI (2.5–10 keV) originates from background sources external to the crystal. These are primarily associated with the cryogenic infrastructure, such as Bremsstrahlung radiation from 210Bi present in the internal Pb shielding, and with environmental backgrounds, mainly due to ambient neutrons and -ray radiation, as previously discussed in Ref. Alloni and others (2025). Dedicated Monte Carlo studies will be performed to assess in more detail the different contributions. To further illustrate this point, Fig. 5 also shows the expected contribution from radioactive bulk contaminations within the operated PbWO4 crystal. This component is evaluated through Monte Carlo simulations of the experimental setup, using as input the results of the high-statistics radioactive screening measurements performed on the same crystal and summarized in Tab. 1.
To evaluate the detection and reconstruction efficiency in the RoI for DM investigations, we injected a comb of simulated mono-energetic pulses directly into the raw data stream and processed them through the full analysis pipeline. For each simulated energy, events that passed the quality selection and were correctly reconstructed were counted as successful detections. The efficiency , dependent on the chosen quality selection tolerance, was then obtained as the fraction of successfully reconstructed events over the total number of injected pulses at each energy value. In Fig. 6, the results of this analysis step are shown. In addition, for the evaluation of the detector baseline resolution, we fit a Gaussian distribution to the simulated lines obtaining: 234 eV, see Fig. 7.
@@ Acknowledgements.
The RES–NOVA collaboration thanks the directors and staff of the Laboratori Nazionali del Gran Sasso for their support, as well as the technical staff involved in the experiment. This work received funding from the European Union’s Horizon Europe programme through the ERC grant ERC–101087295 RES–NOVA. Additional financial support was provided under the National Recovery and Resilience Plan (NRRP) by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU, within the project “Advanced techniques for a next–generation isotopically enriched cryogenic Dark Matter experiment” (2022L2AXP2). This work made use of the Arby software for Geant4–based Monte Carlo simulations, developed within the framework of the Milano–Bicocca activities and currently maintained by O. Cremonesi and S. Pozzi. We gratefully acknowledge the University of Milano–Bicocca and INFN for their continuous support of the collaboration. S. Ghislandi acknowledges support from the U.S. Department of Energy (DOE) Grant No. [DE-SC0011091]. F.A. Danevich and V.I. Tretyak thank the Laboratori Nazionali del Gran Sasso, the Gran Sasso Science Institute, and the RES–NOVA Collaboration for their outstanding support and hospitality during the difficult period following the Russian invasion of Ukraine.References